At an air show, a jet plane has velocity components vx = 675 km/h and vy = 475 km/h at time 4.45 s and vx = 888 km/h and vy = 370 km/h at time 7.37 s .

a) For this time interval, find the x component of the plane's average acceleration. (km/h2 )

b) For this time interval, find the y component of the plane's average acceleration. (km/h2)

c) For this time interval, find the magnitude of its average acceleration. (km/h2)

d) For this time interval, find the direction of its average acceleration. (∘ clockwise from +x axis)
Answer #1

Average acceleration = change in velocity / time interval

------------------------------------------------------------

we need to change time interval from seconds to hours

t = 7.37 - 4.45 = 2.92 sec = 2.92 / 3600 = 8.111e-4 hours

(a) ax = 888 - 675 / 8.111e-4 = 262606.34 km/h2

(b) ay = 370 - 475 / 8.111e-4 = -129453.83 km/h2

(c) |a| = sqrt ( ax2 + ay2)

|a| = 2.927e5 Km/h2

(d) direction = arctan (  129453.83 / 262606.34)

direction = 26.24 degree

or

direction = 333.7 degree ( clockwise from +x axis)